# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ 2.2 Statistical Model for the Site Selection Problem
As in Gechter et al. (2024), the policy maker must choose a strict subset of experimental sites $\mathscr{S}\subset\mathcal{S}_{E}$. [^4] We require $\mathscr{S}$ to be a strict subset of $\mathcal{S}_{E}$ because if we allow the policy maker to experiment in all sites, and there is no cost of experimentation that varies at the site level, then there is no site selection problem. We consider the case in which $\mathscr{S}$ is allowed to equal $\mathcal{S}_{E}$ in Section 6.1. In this paper, we focus on the case in which there is a restriction on the total number of experimental sites that the policy maker can select. That is, there is an integer $k\in\mathbb{N}$, $k<\textrm{card}(\mathcal{S}_{E})$, such that $\mathscr{S}$ must belong to the set
$\mathcal{A}(k):=\{\mathscr{S}\subset\mathcal{S}_{E}\>|\>\textrm{card}(\mathscr{S})\leq k\}.$ (1)
What we have in mind with this restriction is that the policy maker may not have enough administrative resources to run more than $k$ experiments to evaluate the policy of interest. Our notation also allows for the possibility that the policy maker does not want to experiment at all.
If the policy maker decides to experiment in a nonempty set $\mathscr{S}\in\mathcal{A}(k)$ of cardinality $\textrm{card}(\mathscr{S})\leq k$, then she will observe $\textrm{card}(\mathscr{S})$ treatment effect estimates. We collect these estimates in a vector of dimension $\textrm{card}(\mathscr{S})$. In a slight abuse of notation, let $\mathscr{S}_{1}<\mathscr{S}_{2}<\ldots<\mathscr{S}_{\textrm{card}(\mathscr{S})}$ denote the indices of the $\textrm{card}(\mathscr{S})$ experimental sites in $\mathscr{S}$. Letting $\widehat{\mathcal{\tau}}_{s}$ denote the estimated treatment effect in site $s$, we can define the vector
$\widehat{\tau}_{\mathscr{S}}:=(\hat{\tau}_{\mathscr{S}_{1}},\ldots,\hat{\tau}_{\mathscr{S}_{\textrm{card}(\mathscr{S})}})^{\top}.$ (2)
@@
$\tau_{\mathscr{S}}:=(\tau(X_{\mathscr{S}_{1}}),\ldots,\tau(X_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}))^{\top}.$ (3)
@@
$\widehat{\tau}_{\mathscr{S}}\sim\mathcal{N}_{\textrm{card}(\mathscr{S})}\left(\tau_{\mathscr{S}}\>,\>\Sigma_{\mathscr{S}}\right),\textrm{ where }\Sigma_{\mathscr{S}}:=\operatorname{diag}(\sigma_{\mathscr{S}_{1}}^{2},\ldots,\sigma_{\mathscr{S}_{\textrm{card}(\mathscr{S})}}^{2}).$ (4)
Following Gechter et al. (2024), treat $\Sigma_{\mathscr{S}}$ as known.
@@
The normality assumption in (4) is unlikely to hold exactly; however, it is common to assume that treatment effect estimates from randomized controlled trials are asymptotically normal with asymptotic variances that can be estimated consistently. Treating the limiting normal model as an exact finite-sample statistical model then eases exposition and allows us to focus on the core features of the site selection problem. Indeed, working directly with such a limiting model is common in applications of statistical decision theory to econometrics; see Müller (2011) and the references therein for theoretical support and applications in the context of testing problems and Ishihara and Kitagawa (2021), Stoye (2012), or Tetenov (2012) for precedents in closely related work. Gechter et al. (2024) use the same statistical model, but our parameter space has a more specific form as treatment effects are controlled by the Lipschitz function $\mathcal{\tau}$.
[…]
@@
Thus, we can define a treatment rule $T$ as a (measurable) function $T:\mathbb{R}^{\textrm{card}(\mathscr{S})}\rightarrow[0,1]^{\textrm{card}(\mathcal{S}_{P})}$ that maps experimental outcomes to (possibly) randomized policy actions in each of the policy-relevant sites. It will be sometimes convenient to use $T_{s}(\cdot)$ to denote the specific treatment rule for site $s\in\mathcal{S}_{P}$ implied by $T$ and $\mathcal{T}_{\mathscr{S}}$ to denote the set of all treatment rules. Note that we index the treatment rules by the selected experimental sites, $\mathscr{S}$, to be explicit about the fact that the data used to inform policy will vary depending on the choice of $\mathscr{S}$. We call $T\in\mathcal{T}_{\mathscr{S}}$ nonrandomized if for every $s\in\mathcal{S}_{P}$ we have $T_{s}(z)\in\{0,1\}$ for (Lebesgue) almost every $z\in\mathbb{R}^{\textrm{card}(\mathscr{S})}$. Otherwise, we say that the rule is randomized. We will focus on decision rules that belong to the set
$\mathcal{T}_{\mathscr{S}}^{1/2}:=\left\{T\in\mathcal{T}_{\mathscr{S}}\>|\>\textrm{ for any diagonal matrix }\Sigma,\textrm{ and }\forall s\in\mathcal{S}_{P},\mathbb{E}[T_{s}(U)]=1/2,U\sim\mathcal{N}_{\textrm{card}(\mathscr{S})}(\mathbf{0},\Sigma)\right\}.$
These are treatment rules in which the ex-ante probability of implementing the policy is $50\%$ whenever the true treatment effects at the sites experimented on are zero. For the moment, and for the sake of exposition, we assume that there is no cost of experimentation. While this assumption is clearly unrealistic, we later show that the main conclusions of our analysis are robust to adding fixed costs to the objective function of the $k$-median problem.
@@ 2.3 Welfare and Regret
Let #$\#\mathcal{S}_{P}:=\textrm{card}(\mathcal{S}_{P})$. We assume that the welfare of a decision rule $T$, given that sites $\mathscr{S}$ are selected for experimentation, corresponds to the average expected welfare across policy-relevant sites:
@@
where $\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]$ means that the expectation is taken assuming $\widehat{\tau}_{\mathscr{S}}\sim\mathcal{N}_{\textrm{card}(\mathscr{S})}(\tau_{\mathscr{S}},\Sigma_{\mathscr{S}})$, and $\tau(X_{s})$ is the true treatment effect at site $s$.
The regret of policy $(T,\mathscr{S})$ can then be shown to equal
$\mathcal{R}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\left(\mathbf{1}\{\tau(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right).$ (6)
@@
The pair $(T^{*},\mathscr{S}^{*})\in\mathcal{T}_{\mathscr{S}}^{1/2}\times\mathcal{A}(k)$ is minimax-regret (MMR) optimal among all purposive sampling schemes and treatment rules if
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T^{*},\mathscr{S}^{*},\tau)=\inf_{\mathscr{S}\in\mathcal{A}(k),T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau).$ (7)
Remark 1.
In the standard definition of minimax-regret optimality, it is common to allow the decision maker to select randomized decision rules. Definition 1 implies an asymmetric treatment of randomization: While we allow the policy maker to randomize policy implementation choices, we are restricting her to pick the experimental sites in a deterministic fashion. The nonrandom selection of experimental sites—referred to broadly in the literature as purposive sampling (Cook et al., 2002, p. 511)—is common in practice; see the discussion about purposive sampling in Egami and Lee (2024). In Section 6, we discuss the challenges we encountered in trying to allow for the random selection of experimental sites. ∎
@@
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right).$
This suggests that, conceptually, the solutions of the minimax-regret problem can be done in two steps. First, one can analyze the problem of policy implementation given the experimental outcomes at sites $\mathscr{S}$. Then, one can optimize over the sites where to experiment.
This distinction is helpful because the first step is related to the “evidence aggregation” problem analyzed in Ishihara and Kitagawa (2021). As mentioned in the Introduction, we can leverage recent results in Yata (2021) and Montiel Olea et al. (2023) to better understand the minimax-regret solutions to this problem. A key difference from Ishihara and Kitagawa (2021) is that we typically have more than one policy-relevant site. ∎
2.4 Main Assumptions
We make the following assumptions. To begin, treatment effects vary continuously with the site-level covariates:
Assumption 1.
[…]
We also impose a regularity condition on site-level covariates. That is, we assume that all observed covariates are different:
Assumption 2.
$X_{s}\neq X_{s^{\prime}}\ \forall\ s,s^{\prime}\in\mathcal{S}$.
Even if this were not the case in raw data, one would presumably want to induce it by adding site fixed effects. [^5] In section 6.3, we discuss how Assumption 1 can be modified to accommodate other general distance measures, as well as some forms of unobserved treatment heterogeneity (for which Assumption 2 may be dropped).
Finally, for each policy site $s\in\mathcal{S}_{P}$ and every $\mathscr{S}\in\mathcal{A}(k)$, denote by $N_{\mathscr{S}}(s)\in\mathscr{S}$ its nearest neighbor in $\mathscr{S}$ (or the nearest neighbor with the smallest index in case of multiplicity). That is, for every $s\in\mathcal{S}_{P}$:
@@ 2.4 Main Assumptions
3 Main Result
This section presents our main result: The problem of finding the purposive sampling scheme that minimizes the worst-case regret is approximately equal—and sometimes equal—to solving a $k$-median problem.
In order to derive this, first recall the definition of the $k$-median problem.
@@ 3 Main Result
There is no scope for interesting analysis, or indeed for justification of experiments, unless noise in the signal is bounded. Therefore, define $C^{*}:=\underset{\mathscr{S}\in\mathcal{A}(k),s\in\mathcal{S}_{P}\backslash\mathscr{S}}{\max}\left\{\sqrt{\frac{\pi}{2}}\frac{\sigma_{N_{{\mathscr{S}}(s)}}}{\left\|X_{s}-X_{N_{{\mathscr{S}}(s)}}\right\|}\right\}<\infty$. Let $\sigma_{E}$ denote the largest standard deviation among the potential experimental sites. Our main theorem is the following:
@@
$\Bigg{|}\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)-\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)\Bigg{|}$ (10)
@@
Remark 3.
The left-hand expression in (10), i.e., $\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)$, is the worst-case regret of the purposive sampling scheme $\mathscr{S}$ assuming optimality of subsequent treatment choice $T_{\mathscr{S}}$. Theorem 1 thus shows that the (optimized) worst-case regret of any purposive sampling scheme can be uniformly approximated (after scaling by the factor $C/(2\#\mathcal{S}_{P}$)) by the objective function of the k-median problem in Definition 2—which is equivalent to the objective function of the problem in the right-hand side of Equation (9). ∎
Remark 4.
When the candidate sites for experimentation $(\mathcal{S}_{E})$ and the policy-relevant sites $(\mathcal{S}_{P})$ are disjoint, then for any $\mathscr{S}\in\mathcal{A}(k)$:
$\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)=\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$ (12)
Consequently, any solution of the $k$-median problem is a fortiori an exact minimax-regret solution for the site selection problem. This follows from the fact that if a purposive sampling scheme, $\mathscr{S}^{*}$, minimizes the right-hand side of Equation (12), then it will automatically minimize the left-hand side, which, by Remark 2, defines a minimax-regret optimal purposive sampling scheme. ∎
Remark 5.
@@
$\Bigg{|}\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)-\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)\Bigg{|}$ (13)
is at most equal to (11). Thus, the quality of this approximation improves as either the number of policy-relevant sites increases or the potential experiments conducted become more precise ($\sigma_{E}$ becomes smaller). The approximation deteriorates as $k$ increases. This happens because in approximating the worst-case regret of a purposive sampling scheme, we ignored the regret associated with sites selected for experimentation; we further discuss this below when we explain Lemma 1. This means that, as $k$ increases, more sites will be ignored in our approximation and the upper bound will become looser. ∎
Remark 6.
The tight connection to the $k$-median problem relies on the Lipschitz constant to be large enough. In other words, the $k$-median problem only has a decision-theoretic justification when the treatment effect heterogeneity across sites accommodated in the parameter space is substantial. ∎
The proof of Theorem 1 follows from the two Lemmas below, which bound the site selection problem’s minimax-regret value from above and below, respectively.
@@
$\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\leq\;\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}\right)+\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right),$ (14)

### v2

@@ 2.2 Statistical Model for the Site Selection Problem
As in Gechter et al. (2024), the policymaker must choose a strict subset of experimental sites $\mathscr{S}\subset\mathcal{S}_{E}$. [^6] We require $\mathscr{S}$ to be a strict subset of $\mathcal{S}_{E}$ because if we allow the policymaker to experiment in all sites, and there is no cost of experimentation that varies at the site level, then there is no site selection problem. We consider the case in which $\mathscr{S}$ is allowed to equal $\mathcal{S}_{E}$ in Section 6.1. As discussed in the introduction, we focus on the case in which there is a restriction on the total number of experimental sites that the policymaker can select. That is, there is an integer $k\in\mathbb{N}$, $k<\#\mathcal{S}_{E}$, such that $\mathscr{S}$ must belong to the set
$\mathcal{A}(k):=\{\mathscr{S}\subset\mathcal{S}_{E}\>|\>\#\mathscr{S}\leq k\}.$
Our notation also allows for the possibility that the policymaker does not want to experiment at all.
If the policymaker decides to experiment in a nonempty set $\mathscr{S}\in\mathcal{A}(k)$ of cardinality $\mathscr{S}\leq k$, then she will observe $\mathscr{S}$ treatment effect estimates. We collect these estimates in a vector of dimension $\mathscr{S}$. In a slight abuse of notation, let $\mathscr{S}_{1}<\mathscr{S}_{2}<\ldots<\mathscr{S}_{\#\mathscr{S}}$ denote the indices of the $\mathscr{S}$ experimental sites in $\mathscr{S}$. Letting $\widehat{\mathcal{\tau}}_{s}$ denote the estimated treatment effect in site $s$, we can define the vector
$\widehat{\tau}_{\mathscr{S}}:=(\widehat{\tau}_{\mathscr{S}_{1}},\ldots,\widehat{\tau}_{\mathscr{S}_{\#\mathscr{S}}})^{\top}.$
@@
$\tau_{\mathscr{S}}:=(\tau(X_{\mathscr{S}_{1}}),\ldots,\tau(X_{\mathscr{S}_{\#\mathscr{S}}}))^{\top}.$
@@
$\widehat{\tau}_{\mathscr{S}}\sim\mathcal{N}_{\#\mathscr{S}}\left(\tau_{\mathscr{S}}\>,\>\Sigma_{\mathscr{S}}\right),\textrm{ where }\Sigma_{\mathscr{S}}:=\operatorname{diag}(\sigma_{\mathscr{S}_{1}}^{2},\ldots,\sigma_{\mathscr{S}_{\#\mathscr{S}}}^{2}).$ (1)
Following Gechter et al. (2024), we furthermore treat $\Sigma_{\mathscr{S}}$ as known.
@@
The normality assumption in (1) is unlikely to hold exactly; however, it is common to assume that treatment effect estimates from randomized controlled trials are asymptotically normal with asymptotic variances that can be estimated consistently. Treating the limiting normal model as an exact finite-sample statistical model eases exposition and allows us to focus on the core features of the site selection problem. Indeed, working directly with such a limiting model is common in applications of statistical decision theory to econometrics; see Müller (2011) and the references therein for theoretical support and applications in the context of testing problems and Ishihara and Kitagawa (2021), Stoye (2012), or Tetenov (2012) for precedents in closely related work. Gechter et al. (2024) use the same statistical model, but our parameter space has a more specific form as treatment effects are controlled by the Lipschitz function $\mathcal{\tau}$.
After observing $\widehat{\mathcal{\tau}}_{\mathscr{S}}$, the policymaker chooses an action $a_{s}\in[0,1]$ at each policy-relevant site $s\in\mathcal{S}_{P}$. We interpret this action as the proportion of a population in the site that will be randomly assigned to the new policy. Thus, $a_{s}=1$ means that everyone in site $s$ is exposed to the new policy, and $a_{s}=0$ means that the status quo at the site is preserved. Under this interpretation, $a_{s}=.5$ means that 50% of the population at site $s$ will be exposed at random to the new policy; however, the formal development equally applies to either individual or population-level randomization. Our interpretation abstracts from integer issues arising with small populations.
@@
Thus, we can define a treatment rule $T$ as a (measurable) function $T:\mathbb{R}^{\#\mathscr{S}}\rightarrow[0,1]^{\#\mathcal{S}_{P}}$ that maps experimental outcomes to (possibly) randomized policy actions in each of the policy-relevant sites. It will sometimes be convenient to use $T_{s}(\cdot)$ to denote the specific treatment rule for site $s\in\mathcal{S}_{P}$ implied by $T$ and $\mathcal{T}_{\mathscr{S}}$ to denote the set of all treatment rules. Note that we index the treatment rules by the selected experimental sites, $\mathscr{S}$, to be explicit about the fact that the data used to inform policy will vary depending on the choice of $\mathscr{S}$. We call $T\in\mathcal{T}_{\mathscr{S}}$ nonrandomized if for every $s\in\mathcal{S}_{P}$ we have $T_{s}(z)\in\{0,1\}$ for (Lebesgue) almost every $z\in\mathbb{R}^{\#\mathscr{S}}$. Otherwise, we say that the rule is randomized. For the moment, and for the sake of exposition, we assume that there is no cost of experimentation. While this assumption is clearly unrealistic, we later show that the main conclusions of our analysis are robust to adding fixed costs to the objective function of the $k$-median problem.
Our setup only allows for treatment effect heterogeneity as a function of site-level covariates. This is partly motivated by the fact that detailed individual-level data may not be available for all sites in empirical applications. If individual-level covariates were available, and if there were no heterogeneity of CATEs (defined via individual-level characteristics) across sites, one could estimate CATEs with data from experimental sites and then reweight them to derive average treatment effects for policy-relevant sites; see, for example, the discussion in List (2024, p. 493). In this case, many state-of-the-art methods for estimating ATE with experimental data could be employed. However, one might still be concerned about external validity, namely that those CATEs vary at the site level. A pure reweighting method would then not work, and some extrapolation would still be required. Our approach already offers a practical solution to deal with the individual-level covariates, at least in the case in which there is a single policy-relevant site: One can estimate site-specific CATEs and take $\widehat{\tau}_{s}$ to be the average of these CATEs over the distribution of covariates in the policy-relevant site. Then we are back to the original set-up of our problem, in which we need to decide how to aggregate these transported estimators to make policy choices in the site of interest. Since the payoff relevant parameter continues to be the site-level treatment effect, we think it is reasonable to model treatment effect heterogeneity as a function of the site-level characteristics.
@@ 2.3 Welfare and Regret
We assume that the welfare of a decision rule $T$, given that sites $\mathscr{S}$ are selected for experimentation, corresponds to the average expected welfare across policy-relevant sites:
@@
where $\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]$ means that the expectation is taken assuming $\widehat{\tau}_{\mathscr{S}}\sim\mathcal{N}_{\#\mathscr{S}}(\tau_{\mathscr{S}},\Sigma_{\mathscr{S}})$, and $\tau(X_{s})$ is the true treatment effect at site $s$.
The regret of policy $(T,\mathscr{S})$ equals
$\mathcal{R}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\bigl{(}\mathbf{1}\{\tau(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\bigr{)}.$ (3)
@@
The pair $(T^{*},\mathscr{S}^{*})\in\mathcal{T}_{\mathscr{S}}\times\mathcal{A}(k)$ is minimax-regret (MMR) optimal among all purposive sampling schemes and treatment rules if
$\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T^{*},\mathscr{S}^{*},\tau)=\inf_{\mathscr{S}\in\mathcal{A}(k),T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau).$ (4)
In the standard definition of MMR optimality, the decision maker may select randomized decision rules. Definition 1 implies an asymmetric treatment of randomization: While we allow the policymaker to randomize policy implementation choices, we are restricting her to pick the experimental sites in a deterministic fashion. In Section 6, we discuss challenges we encountered in trying to allow for the random selection of experimental sites.
@@
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right).$
This suggests that, conceptually, the MMR problem can be solved in two steps. First, analyze the problem of policy implementation given the experimental outcomes at sites $\mathscr{S}$; then, optimize over the sites where to experiment.
This distinction is helpful because the first step is related to Ishihara and Kitagawa’s (2021 “evidence aggregation” problem, a key difference being that we typically have more than one policy-relevant site. ∎
3 Main Result
This section presents our main result: Finding the MMR purposive sampling scheme is approximately, and sometimes exactly, equal to solving a $k$-median problem. To develop this, for each policy site $s\in\mathcal{S}_{P}$ and every $\mathscr{S}\in\mathcal{A}(k)$, denote by $N_{\mathscr{S}}(s)\in\mathscr{S}$ its nearest neighbor in $\mathscr{S}$ (or the nearest neighbor with the smallest index in case of multiplicity). That is, for every $s\in\mathcal{S}_{P}$:
@@ 3 Main Result
Recall the following definition:
@@
Our main result depends on treatment effect heterogeneity being possibly substantial. To formalize this, define $C^{*}:=\underset{\mathscr{S}\in\mathcal{A}(k),s\in\mathcal{S}_{P}\backslash\mathscr{S}}{\max}\left\{\sqrt{\frac{\pi}{2}}\frac{\sigma_{N_{{\mathscr{S}}(s)}}}{\left\|X_{s}-X_{N_{{\mathscr{S}}(s)}}\right\|}\right\}<\infty$. Let $\sigma_{E}$ denote the largest standard deviation among the potential experimental sites. We then have:
@@
$\Bigg{|}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)-\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)\Bigg{|}$ (7)
@@
The left-hand expression in (7), i.e., $\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)$, is the worst-case regret of the purposive sampling scheme $\mathscr{S}$ assuming optimality of subsequent treatment choices. Theorem 1 thus shows that the (optimized) worst-case regret of any purposive sampling scheme can be uniformly approximated (after scaling by the factor $C/(2\#\mathcal{S}_{P}$)) by the objective function of the k-median problem in Definition 2, which in turn is the objective function on the right-hand side of Equation (6). Note that our Assumption 1 implies that the treatment effect heterogeneity between site $s$ and $s^{\prime}$ is governed by the product of $C$ and $\|X_{s}-X_{s^{\prime}}\|$. Given $\mathscr{S}\in\mathcal{A}(k)$, a smaller value of $\left\|X_{s}-X_{N_{{\mathscr{S}}(s)}}\right\|$ for $s\in\mathcal{S}_{P}$ indicates more similarity between policy and experimental sites, i.e., less treatment heterogeneity with $C$ fixed. Consequently, a larger $C^{*}$ is needed to allow for overall treatment heterogeneity to be substantial.∎
Under conditions of Theorem 1, when the candidate sites for experimentation $(\mathcal{S}_{E})$ and the policy-relevant sites $(\mathcal{S}_{P})$ are disjoint, then for any $\mathscr{S}\in\mathcal{A}(k)$:
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)=\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|.$ (9)
Consequently, any solution of the $k$-median problem is then an exact minimax-regret solution for the site selection problem. This follows from the fact that if a purposive sampling scheme, $\mathscr{S}^{*}$, minimizes the right-hand side of Equation (9), then it will also minimize the left-hand side, which, by Remark 1, defines a minimax-regret optimal purposive sampling scheme. ∎
@@
$\Bigg{|}\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)-\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)\Bigg{|}$
is at most equal to (8). Thus, the quality of this approximation improves as either the number of policy-relevant sites increases or the potential experiments conducted become more precise ($\sigma_{E}$ becomes smaller). The approximation deteriorates as $k$ increases. This happens because in approximating the worst-case regret of a purposive sampling scheme, we ignored the regret associated with sites selected for experimentation; see also further discussion of Lemma 1 below. This means that, as $k$ increases, more sites will be ignored in our approximation and the upper bound will become looser. ∎
Theorem 1 follows from the following two lemmas, which bound the site selection problem’s MMR value from above and below.
@@
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)~\leq~\;\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}~+~\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|,$